# Incorporating Task Progress Knowledge for Subgoal Generation in Robotic Manipulation through Image Edits

### v1

@@ 4.4 Impact of Dataset Size
Our real robot experiments show that we can use much fewer demonstrations (1.4k trajectories recorded over 7 hours) than prior works, e.g. 60k trajectories used in the baseline method SuSIE, to generate subgoal images. To understand the impact of training set size, we further train the models with various dataset sizes on CALVIN. We sampled 75%, 50%, and 25% of the data (i.e., an average of 97, 65, and 33 trajectories for each task) to train all methods to assess how performance changes with reduced data. Tab. 4 shows that even with 25% of data, our success rate remain relatively high. The example subgoal generations can be found in Appendix A.
@@
Table 4: Success rates (%) on the CALVIN dataset with different amounts of training data. TaKSIE has a good success rate even when trained on only 25% of the available data.
@@ 4.5 Generalization to Unseen Scenarios
TaKSIE (Ours) 54.31 43.51
SuSIE 49.64 41.32
@@
Unseen Environment: We selected the same 9 tasks as in Unseen tasks to train TaKSIE in three CALVIN environments (A, B, C) and tested the model on the same tasks but in an unseen environment D. Appendix B shows the train and test environments with the example generations.
Tab. 5 shows that the success rate drops in both cases but ours still outperforms baselines.
@@ 4.7 Failure Cases
We observe three types of failure cases. First, the model may generate incorrect subgoals, for example, it generates pushing a block to the right in the task of pushing it to the left. This indicates that language embedding does not provide sufficient spatial information. Using a more powerful language encoder can help in the future. Second, the model may generate low-quality images, e.g., blurry or distorted regions. Having bad image quality affects the performance of the goal-conditioned policy as it cannot identify the correct visual goals. In TaKSIE, we regenerate subgoal images when the robot does not achieve the subgoal in $\lambda$ time steps which mitigates the impact of the low-quality images. Third, the policy may fail to reach the subgoal when the generated subgoal is good. This indicates that the improvement in the goal-conditioned policy is needed.
@@ 5 Conclusion & Future Work
We have demonstrated that it is possible to incorporate task progress to generate better visual subgoals to guide the rollout of policies. Instead of using a fixed interval, we showed that the task progress implicitly encoded in the time-contrastive visual representations can help select ground-truth subgoals. Together with a progress encoder and evaluator, TaKSIE can boost the performance of robotic manipulation tasks in both simulation and real-world environments. While our design separates subgoal generation and plan execution into two stages, we may consider cotrain the subgoal generator and the low-level policy so both models can be aware of each other’s capability to align representations better. Our current task is fully observable and mostly atomic and sequential. Future work can extend to mobile manipulation and complex task structures. In these scenarios, handling partial observation, changing viewpoints, and dependency of subgoals will be important.
@@ Appendix B Example of subgoal generation on unseen tasks and environment
Appendix C Ablation: Slope Values
The slope parameters can affect the selection of subgoals. We experiment with different $\delta_{1}$ and $\delta_{2}$ values in 9 selected tasks in each category of the CALVIN validation set. Tab. 6 shows that smaller $\delta_{1}$ and larger $\delta_{2}$ lead to capture more subgoals. Overall, the number of our selected subgoals is less than SuSIE which selects 3 subgoals on average. Compared to SuSIE’s success rate (79.73% in Tab. 4), most of our slope values can perform better, indicating that our selected subgoals can still guide the policy. However, our experiments show no direct relationship between the number of selected subgoals and the success rate as the performance depends on the informativeness of the selected subgoals than just numbers.

### v2

@@ 4.4 Impact of Dataset Size
Our real robot experiments show that we can use fewer demonstrations (1.4k trajectories recorded over 7 hours) than prior works, e.g. 60k trajectories used in SuSIE, to generate subgoal images. To understand the impact of training set size, we further train the models with various dataset sizes on CALVIN. We sampled 75%, 50%, and 25% of the data (i.e., an average of 97, 65, and 33 trajectories for each task) to train all methods to assess how performance changes with reduced data. Tab. 4 shows that even with 25% of data, our success rate remains relatively high. The example subgoal generations can be found in Appx. A.
@@
Table 4: Success rates (%) on the CALVIN dataset with different amounts of training data. TaKSIE has a relatively good success rate even when trained on only 25% of the available data.
@@ 4.5 Generalization to Unseen Scenarios
TaKSIE (Ours) 54.31 73.74
SuSIE 49.64 71.24
@@
Unseen Environment: We utilized all 34 tasks as in unseen tasks to train TaKSIE in three CALVIN environments (A, B, C) and tested the model on the same tasks but in an unseen env. D (see Appx. B for example environments and generations). Tab. 5 shows that the success rate drops in both cases but ours still outperforms baselines.
@@ 4.7 Failure Cases
We observe three types of failure cases (see examples in Appx. C). First, the model may generate incorrect subgoals, for example, it generates pushing a block to the right in the task of pushing it to the left. This indicates that language embedding does not provide sufficient spatial information. Using a more powerful language encoder can help in the future. Second, the model may generate low-quality images, e.g., blurry or distorted regions. Having bad image quality affects the performance of the goal-conditioned policy as it cannot identify the correct visual goals. In TaKSIE, we regenerate subgoal images when the robot does not achieve the subgoal in $\lambda$ time steps which mitigates the impact of the low-quality images. Third, the policy may fail to reach the subgoal when the generated subgoal is good. This indicates that improvement in the policy is needed.
@@ 5 Conclusion & Future Work
We have demonstrated that it is possible to incorporate task progress to generate better visual subgoals to guide the rollout of policies. Instead of using a fixed interval, we showed that the task progress implicitly encoded in the time-contrastive visual representations can help select ground-truth subgoals. Together with a progress encoder and evaluator, TaKSIE can boost the performance of robotic manipulation tasks in simulation and the real world. While our design separates subgoal generation and policy execution into two stages, we may consider cotrain them so both models can be aware of each other’s capability to align representations better. Our current task is fully observable and mostly atomic and sequential. Future work can extend to mobile manipulation and complex task structures. In these scenarios, handling partial observation, changing viewpoints, and dependency of subgoals will be important.
6 Acknowledgment
The authors acknowledge Research Computing at the University of Virginia for providing the computational resources and technical support that made the results in this work possible.
@@ Appendix B Example of subgoal generation on unseen tasks and environment
Appendix C Example of failures cases
Figure 10: Three common failure cases: (1) incorrect subgoals, (2) low-quality images, and (3) failure to reach valid subgoals.
Appendix D Comparsion of different image-conditioned diffusion model
Figure 11: Comparison of subgoal generation using different image-conditioned diffusion models.
In our tests, both InstructPix2Pix and ControlNet demonstrate similar performance in seen environments. However, for unseen environments, we observe that InstructPix2Pix tends to perform better at following the conditioned environment. Fig. 11 provides an example comparison: both ControlNet and InstructPix2Pix are trained on environments A, B, and C from CALVIN and tested on the unseen environment D. InstructPix2Pix generates subgoals that appear more aligned with the conditions of environment D, while ControlNet is more aligned with environment C, showing less effective generalization to unseen environments. This observation suggests that InstructPix2Pix may have an advantage in generalizing to new scenarios.
Appendix E Ablation: Slope Values
The slope parameters can affect the selection of subgoals. We experiment with different $\delta_{1}$ and $\delta_{2}$ values in all tasks of the CALVIN validation set. Tab. 6 shows that smaller $\delta_{1}$ and larger $\delta_{2}$ lead to capture more subgoals. Overall, the number of our selected subgoals is less than SuSIE which selects 3 subgoals on average. Compared to SuSIE’s success rate (79.73% in Tab. 4), most of our slope values can perform better, indicating that our selected subgoals can still guide the policy. However, our experiments show no direct relationship between the number of selected subgoals and the success rate as the performance depends on the informativeness of the selected subgoals rather than the number of subgoals.